OpenStudy (anonymous):

what is the best integer estimate for the square root 21

OpenStudy (kuoministers):

lol... Integer esimate? well sqrt is kinda halfway between 4 and 5.

OpenStudy (kuoministers):

@uri any idea what is a integer estimate?

OpenStudy (anonymous):

my teacher gave me a worksheet and im trying to do it at home the instructions are find the best integer estimate for each square root then check your estimate with a calculatetor the first number is 21

OpenStudy (kuoministers):

cool what grade is this?

OpenStudy (anonymous):

im in 11th but im bad at math so this should be pre algebra so like 9th or 10th

OpenStudy (kuoministers):

well think this... 3 times 3 is 9 4 times 4 is 16 5 times 5 is 25 so sqrt 21 would be closest to 5

OpenStudy (anonymous):

so it basicly its a number times the same number that equals the number like 4 times 4 is 16

OpenStudy (kuoministers):

yea

OpenStudy (kuoministers):

and 4 and 5 are closest to 21... 5 is closer so that should be the answer

OpenStudy (anonymous):

since this is estimate their might not be exact numbers so 5 to the second power is closest right?

OpenStudy (kuoministers):

yes

OpenStudy (anonymous):

k thxs

OpenStudy (anonymous):

do i put just 5 or five to the second power

OpenStudy (kuoministers):

5

OpenStudy (kuoministers):

if you think my answer helped you please click the best responce button

OpenStudy (anonymous):

i did you helped alot

OpenStudy (kuoministers):

\[\Huge{\color{Purple}{\\{\ \boxed{\mathbb{\text{}{}}{\color{green}{\color{blue}T{\color{purple}H{\color{pink}A{\color{red}NK{\color{orange}S}}}}}}}}}} \]

OpenStudy (kuoministers):

you might wanna work on some english too :D just joking

OpenStudy (anonymous):

no problem thanks for the help i have 7 more to do now ha

OpenStudy (unklerhaukus):

4.5x4.5=4.5x4+4.5x0.5 = 18+2.25 = 20.25

OpenStudy (anonymous):

hu?

OpenStudy (kuoministers):

a integer must be a whole number and cannot have a decimal place.. I think you confused him haha

OpenStudy (anonymous):

well i have a bunch more to do so if anyone isnt busy i could alwayse use help but i am a slow learner ha